# On the transitivity of Anosov diffeomorphisms

### v2

@@ 5. Proof of Theorem B
Theorem B is a particular instance of more general results that we shall prove in this section. The next result comprehensibly contains item (a) of Theorem B.
Theorem 5.1.
@@
$(i)\Rightarrow(2i)$ This is [14][Lemma 2.4].
@@
we get $\overline{\operatorname{Per}(T^{k})}=\Omega(T^{k})$. Thus, we may apply Lemma 3.1 to $T^{k}$ and this way conclude that $T^{k}$ is chain-transitive. This completes the proof of the lemma. ∎
Let us resume the proof of the implication $(3i)\Rightarrow(4i)$. Suppose that the homeomorphism $T$ has the $su-$intersecting property in $\operatorname{Per}(T)$. By Lemma 5.2, $T^{k}$ is chain-transitive for every positive integer $k$. Moreover, as $T$ has the shadowing property, $T^{k}$ has this property as well for every positive integer $k$ (cf. [6][Theorem 2.3.3]). In addition, a chain-transitive homeomorphism with the shadowing property is topologically transitive. So, $T^{k}$ is topologically transitive for every positive integer $k$. Therefore, by [30][Theorem 45], $T$ has the specification property.
$(4i)\Rightarrow(i)$ This is [14][Lemma 2.2]. ∎
@@
$(3i)\Rightarrow(4i)$ This is [59][Lemma 2.2].
$(4i)\Rightarrow(i)$ Since $T$ is expansive and has the shadowing property, one has $\overline{\operatorname{Per}(T)}=\Omega(T)$ according to [6, Theorem 3.1.8]. So, if the homeomorphism $T$ has the barycenter property in $\operatorname{Per}(T)$ then, by Theorem A, $T$ is topologically transitive. Since $T$ is topologically transitive, has the shadowing property and $X$ is connected, then $T$ is topologically mixing (cf. [15][Proposition 3]). Therefore, $T$ has the specification property (cf. [30][Theorem 45]). Hence, by Theorem 5.1, $T$ has the $su-$intersecting property in $X$. ∎
@@ 6. Proof of Theorem C
Let $M$ be a smooth closed Riemannian manifold with dimension $\geqslant 2$ and $f\colon M\to M$ be a $C^{1}$ diffeomorphism. Recall (from the Introduction) that if $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then $f$ has the barycenter property in $\operatorname{Per}(f)$; furthermore, if $\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$, then $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$ if and only if $f$ has the barycenter property in $\operatorname{Per}(f)$. Thus, within the family of $\Omega-$stable $C^{1}$ diffeomorphims we may use one or the other property. Theorem C is a particular case of the more general comparison between several dynamical properties we will now establish.
Theorem 6.1.
Let $M$ be a smooth closed Riemannian manifold, with dimension at least $2$, and $f\colon M\to M$ be an $\Omega-$stable $C^{1}$ diffeomorphism. Then the following statements are equivalent:
  • $(1)$
    $f$ has the two-sided limit shadowing property.
  • $(2)$
    $f$ has the $su-$intersecting property in $M$.
  • $(3)$
    $f$ has the $su-$intersecting property in $\operatorname{Per}(f)$.
  • $(4)$
    $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$.
  • $(5)$
    There exists $A\subseteq M$ such that $f(A)=A$, $\Omega(f)\subseteq\overline{A}$ and $f$ has the barycenter property in $A$.
  • $(6)$
    $f$ is a topologically transitive Anosov diffeomorphism.
  • $(7)$
    $f$ has the specification property.
  • $(8)$
    $f$ has the almost specification property.
  • $(9)$
    $f$ has the asymptotic average shadowing property.
  • $(10)$
    $f$ has the average shadowing property.
  • $(11)$
    $f$ is chain-transitive.
  • $(12)$
    $\mathcal{P}_{f}^{erg}(M)$ is dense in $\mathcal{P}_{f}(M)$.
  • $(13)$
    $f$ has the gluing orbit property.
Proof.
We will use the following schedule:
@@
$(1)$Rightarrow $(2)$Rightarrow $(3)$Rightarrow $(4)$Rightarrow $(5)$Rightarrow $\,(6)$Rightarrow $(7)$Rightarrow $(8)$Rightarrow $(9)$Rightarrow $(10)$Rightarrow $(11)$
\,\,\,$\Downarrow$
$(12)$Rightarrow $(11)$Rightarrow $(1)$
$(1)\Rightarrow(2)$ This is [14][Lemma 2.4].
@@
$(3)\Rightarrow(4)$ This is also obvious since $W^{s}_{f}(x)\cap W^{u}_{f}(y)\subseteq W^{s}_{f}\big{(}\mathcal{O}_{f}(x)\big{)}\cap W^{u}_{f}\big{(}\mathcal{O}_{f}(y)\big{)}$.
$(4)\Rightarrow(5)$ Suppose that $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. According to [59][Lemma 2.2], $f$ has the barycenter property in $\operatorname{Per}(f)$. Since $f$ is Axiom $A$, one has $\Omega(f)\,=\,\overline{\operatorname{Per}(f)}$.
@@
As, in addition, $f$ has the barycenter property in $A$, $f(A)=A$ and $\Omega(f)\subseteq\overline{A}$, we already know from Lemma 4.1 that $f$ is chain-transitive. So $CR(f)=M$. Consequently, $\Omega(f)=M$. Since $\Omega(f)$ is hyperbolic, we conclude that $M$ is hyperbolic, that is, $f$ is an Anosov diffeomorphism. Thus, $f$ has the shadowing property. Besides, as already mentioned, a chain-transitive homeomorphism with the shadowing property is topologically transitive. Hence $f$ is a topologically transitive Anosov diffeomorphism.
$(6)\Rightarrow(7)$ Suppose that $f$ is a topologically transitive Anosov diffeomorphism. So $f$ is an expansive topologically transitive diffeomorphism that has the shadowing property. Since $M$ is connected, then $f$ is topologically mixing [15][Proposition  3]. Therefore, by [30, Theorem 45], $f$ has the specification property.
$(7)\Rightarrow(8)$ This is a consequence of [47][Proposition 2.1]. See also the remarks on [61, page 4] and [29, page 249].
$(8)\Rightarrow(9)$ This is [29][Theorem 3.5].
$(9)\Rightarrow(10)$ This is [29][Theorem 3.7].
$(10)\Rightarrow(11)$ This is [29][Lemma 3.1].
$(11)\Rightarrow(1)$ Since $f$ is $\Omega-$stable and, in addition, chain-transitive, we conclude as previously that $\Omega(f)=\overline{\operatorname{Per}(f)}=CR(f)=M$ and $\Omega(f)$ is hyperbolic. So $f$ is a chain-transitive diffeomorphism satisfying the shadowing property, and therefore topologically transitive. By [14][Corollary 2.3], the topologically transitive Anosov diffeomorphism $f$ has the two-sided limit shadowing property.
@@
As explained before, by the $\Omega-$stability of $f$, one has $\Omega(f)=\overline{\operatorname{Per}(f)}=CR(f)$ and $CR(f)$ is hyperbolic. Besides, by definition, every periodic point of $f$ is an element of the measure center $\mathcal{M}(M,f)$. Moreover, due to [16][Lemma 5.2], one has
@@
Now, with the equalities $\overline{\mathcal{P}_{f}^{erg}(M)}=\mathcal{P}_{f}(M)$ and $\mathcal{M}(M,f)=CR(f)$, we deduce from [16][Lemma 5.8] that $f$ is topologically transitive. In particular, $f$ is chain-transitive.
@@
Scholium 6.3.
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points. If $T$ is topologically transitive, then for every $\xi>0$ there exists a positive integer $N=N(\xi)$ such that for any $a,b\in X$ there is a $\xi-$chain from $a$ to $b$ with length $p$, where $2\leq p\leq N$.
Proof.
Recall that one-sided topological transitivity is equivalent to topological transitivity whenever the compact metric space does not contain isolated points (cf. [4][Theorem 1.4]). So, $T$ is one-sided topologically transitive. Consider an open finite subcover of $X$, given by $\{U_{1}\cdots,U_{s}\}$, such that $\operatorname{diam}(U_{i})<\xi$ for every $i\in\{1,\cdots,s\}$. Thus, by the one-sided transitivity, for each pair $(i,j)\in\{1,\cdots,s\}\times\{1,\cdots,s\}$ there exists $t_{(i,j)}\in\mathbb{N}$ such that $T^{-t_{(i,j)}}(U_{i})\cap U_{j}\neq\emptyset$. Define
[…]
We claim that for any $a,b\in X$ there is a $\delta-$chain from $a$ to $b$ of length $p$ with $2\leq p\leq N$. Indeed, since $X=\bigcup_{i=1}^{s}U_{i}$, there are $r,m$ such that $T(a)\in U_{r}$ and $b\in U_{m}$. Using that $T^{-t_{(m,r)}}(U_{m})\cap U_{r}\neq\emptyset$, take $c\in U_{r}$ such that $T^{t_{(m,r)}}(c)\in U_{m}$. Then, define a $\xi-$pseudo-orbit as follows:
[…]
Note that the length of this chain is $p=t_{(m,r)}+2$, hence $2\leq p\leq N$. ∎
By Scholium 6.3, there exists $N=N(\delta)$ such that we may construct:
@@
  • a $\delta-$chain $(x_{k-1,k})_{0}^{N-1}$ from $T^{n_{k-1}}(y_{k-1})$ to $y_{k}$ with length $p_{(k-1)}$ where $2\leq p_{(k-1)}\leq N$ with
    $(x_{k-1,k})_{0}\,=\,T^{n_{k-1}}(y_{k-1})\quad\quad\text{and}\quad\quad(x_{k-1,k})_{p_{(k-1)}-1}\,=\,y_{k}.$
@@
In [16][Proposition 5.7], the authors showed that, assuming the shadowing property, then the topological transitivity is equivalent to the denseness of the ergodic measures in the space of invariant probability measures. So, the equivalence $(5)\Leftrightarrow(11)$ is an $\Omega-$stable counterpart of that result.
@@
The proof of Theorem 6.1 also ensures, without the assumption of $\Omega-$stability, that
two-sided limit shadowing $\Rightarrow$ $su-$intersecting property in $M$ (6.1)
$\Rightarrow$ $su-$intersecting property in $\operatorname{Per}_{h}(f)$
$\Rightarrow$ weak suintersecting property in $\quad\text{weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$}.$
Moreover (see [53][Lemma 2.1]),
the specification propertyweak suintersecting property in $\text{the specification property}\quad\Rightarrow\quad\text{weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$}.$
@@ 7. Applications
Let $M$ be a smooth closed Riemannian manifold and ${\rm Diff}^{1}(M)$ stand for the set of all $C^{1}$ diffeomorphisms of $M$, endowed with the $C^{1}$ topology. On the next pages we discuss a few consequences of our main results. They comprise describing both the $C^{1}$ interior and a $C^{1}$ generic subset of the space of diffeomorphisms which satisfy the barycenter property in the set of periodic points; and showing that, $C^{1}$ generically in the space of volume-preserving diffeomorphisms, satisfying the barycenter property in the set of hyperbolic periodic points is equivalent to being a transitive Anosov diffeomorphism.
@@ 7.1. Application 1
There is a $C^{1}$ Baire generic set $\mathcal{R}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}$ and there exists $A\subseteq M$ such that $f(A)=A$, $\Omega(f)\subseteq\overline{A}$ and $f$ has the barycenter property in $A$, then $f$ is topologically mixing.
@@
Recall that, from [10][Corollaire 1.2], there exists a Baire generic set $\mathcal{R}_{1}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}_{1}$, then $\Omega(f)=CR(f)$. Moreover, as we are assuming that $M$ is connected, by [10][Corollaire 1.3] there exists a Baire generic set $\mathcal{R}_{2}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}_{2}$ and $M=CR(f)$, then $f$ is topologically transitive. From [3][Theorem 2], there exists a Baire generic set $\mathcal{R}_{3}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}_{3}$ and $f$ is topogically transitive, then $f$ is topogically mixing. Take $f\in\mathcal{R}_{1}\cap\mathcal{R}_{2}\cap\mathcal{R}_{3}$ and suppose that, for some $A\subseteq M$, we have $f(A)=A$, $\Omega(f)\subseteq\overline{A}$ and $f$ has the barycenter property in $A$. Thus, by Lemma 4.1, $f$ is chain-transitive, and so $CR(f)=M$. Thus, $f$ is topologically transitive. Therefore, $f$ is topologically mixing. ∎
@@ 7.2. Application 2
Since an Anosov diffeomorphism $f\colon M\to M$ is structurally stable (so the intersections of every stable manifold with every unstable manifold are transversal) and, by Theorem C, a topologically transitive Anosov diffeomorphism has the weak $su-$intersecting property in $\operatorname{Per}(f)$, we conclude that, under the transitivity assumption, every two periodic points of $f$ are homoclinically related. Consequently, Theorem C implies that:
@@
Adapting the proof of [14][Corollary 3.2] we obtain:
@@
Let $p,q$ be two distinct periodic points of $f$. As $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, the sets $W^{s}_{f}(\mathcal{O}_{f}(p))\cap W^{u}_{f}(\mathcal{O}_{f}(q))$ and $W^{u}_{f}(\mathcal{O}_{f}(p))\cap W^{s}_{f}(\mathcal{O}_{f}(q))$ are nonempty. Taking into account that $f\in\mathcal{KS}$, we know that these intersections are transversal. Therefore, $p$ and $q$ are homoclinically related, and so they have the same index. ∎
@@
$\mathcal{PCI}\,=\,\big{\{}f\in{\rm Diff}^{1}(M)\colon\,\forall\,p,q\in\operatorname{Per}_{h}(f)\quad\operatorname{ind}(p)=\operatorname{ind}(q)\big{\}}.$
@@
[53][Lemma 2.4] Let $G\in{\rm Diff}^{1}(M)$ and $P$ be a non-hyperbolic periodic point of $G$. Then, for every $C^{1}$ neighborhood $\mathcal{V}$ of $G$ there are $H\in\mathcal{V}$ and $P_{1},P_{2}\in\operatorname{Per}_{h}(H)$ such that $\operatorname{ind}(P_{1})\neq\operatorname{ind}(P_{2})$.
By Lemma 7.5, there exist $h\in\mathcal{U}$ and two distinct hyperbolic periodic points $p_{1},p_{2}$ of $h$ with different indices. Yet, this contradicts the fact that $h\in\mathcal{PCI}$. Therefore, $f$ is a star diffeomorphism, and so $\Omega-$stable.
Consequently:
@@
Suppose that $f$ is chain-transitive and $f$ is in $C^{1}$ interior of $\mathcal{PCI}$. By Proposition 7.4, the diffeomorphism $f$ is $\Omega-$stable. Since $f$ is also chain-transitive, by Theorem 6.1 $f$ is a topologically transitive Anosov diffeomorphism.
The converse assertion is clear, due to the openess of the set of topologically transitive Anosov diffeomorphisms and Corollary 7.2. ∎
The next proposition completes the previous information on a $C^{1}$ generic subset of ${\rm Diff}^{1}(M)$.
@@
[14][Lemma 3.3] There exists an $C^{1}$ open and dense set $\mathcal{I}$ of ${\rm Diff}^{1}(M)$ such that every $f\in\mathcal{I}$ has this property: if for any $C^{1}$ neighborhood $\mathcal{U}$ of $f$ some $g\in\mathcal{U}$ has two hyperbolic periodic points with different indices, then $f$ also has two hyperbolic periodic points with different indices.
@@
If $f\in\mathcal{R}\,\cap\,\mathcal{PCI}$ and is chain-transitive, then $f$ is a topologically transitive Anosov diffeomorphisms.
@@
Take a chain-transitive $f\in\mathcal{R}\,\cap\,\mathcal{PCI}$. By Proposition 7.7, $f$ is $\Omega-$stable. Then, by Theorem 6.1, $f$ is a topologically transitive Anosov diffeomorphism. ∎

### v3

@@ 5. Proof of Theorem B
Theorem B is a particular instance of more general results that we shall prove in this section. The next one comprehensibly contains item (a) of Theorem B.
Proposition 5.1.
@@
$(i)\Rightarrow(2i)$ This is [14, Lemma 2.4].
@@
we get $\overline{\operatorname{Per}(T^{k})}=\Omega(T^{k})$. Thus, we may apply Lemma 3.1 to $T^{k}$ and this way conclude that $T^{k}$ is chain-transitive. ∎
Let us resume the proof of the implication $(3i)\Rightarrow(4i)$. Suppose that the homeomorphism $T$ has the $su-$intersecting property in $\operatorname{Per}(T)$. By Lemma 5.2, $T^{k}$ is chain-transitive for every positive integer $k$. Moreover, as $T$ has the shadowing property, $T^{k}$ has this property as well for every positive integer $k$ (cf. [6, Theorem 2.3.3]). In addition, a chain-transitive homeomorphism with the shadowing property is topologically transitive. So, $T^{k}$ is topologically transitive for every positive integer $k$. Therefore, by [30, Theorem 45], $T$ has the specification property.
$(4i)\Rightarrow(i)$ This is [14, Lemma 2.2]. ∎
@@
$(3i)\Rightarrow(4i)$ This is [58, Lemma 2.2].
$(4i)\Rightarrow(i)$ Since $T$ is expansive and has the shadowing property, then $\overline{\operatorname{Per}(T)}=\Omega(T)$ by [6, Theorem 3.1.8]. So, if the homeomorphism $T$ has the barycenter property in $\operatorname{Per}(T)$ then, by Theorem A, $T$ is topologically transitive. Since $T$ is topologically transitive, has the shadowing property and $X$ is connected, then $T$ is topologically mixing (cf. [15, Proposition 3]). Therefore, $T$ has the specification property (cf. [30, Theorem 45]). Hence, by Proposition 5.1, $T$ has the $su-$intersecting property in $X$. ∎
@@ 6. Proof of Theorem C
We will use the following scheme of implications:
@@
$(1)$Rightarrow $(2)$Rightarrow $(3)$Rightarrow $(4)$Rightarrow $(5)$Rightarrow $\,(6)$Rightarrow $(7)$Rightarrow $(8)$Rightarrow $(9)$Rightarrow $(10)$
\,\,\,$\Downarrow$ \,\,\,$\Downarrow$
$(12)$Rightarrow $(11)$Rightarrow $(1)$ $(11)$
$(1)\Rightarrow(2)$ This is [14, Lemma 2.4].
@@
$(3)\Rightarrow(4)$ This is also immediate since $W^{s}_{f}(x)\cap W^{u}_{f}(y)\subseteq W^{s}_{f}\big{(}\mathcal{O}_{f}(x)\big{)}\cap W^{u}_{f}\big{(}\mathcal{O}_{f}(y)\big{)}$.
$(4)\Rightarrow(5)$ Suppose that $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. According to [58, Lemma 2.2], $f$ has the barycenter property in $\operatorname{Per}(f)$. Since $f$ is Axiom $A$, one has $\Omega(f)\,=\,\overline{\operatorname{Per}(f)}$.
@@
As, in addition, $f$ has the barycenter property in $A$, $f(A)=A$ and $\overline{A}=\Omega(f)$, we already know from Lemma 4.1 that $f$ is chain-transitive. So $CR(f)=M$. Consequently, $\Omega(f)=M$. Since $\Omega(f)$ is hyperbolic, we conclude that $M$ is hyperbolic, that is, $f$ is an Anosov diffeomorphism. Thus, $f$ has the shadowing property. Besides, as already mentioned, a chain-transitive homeomorphism with the shadowing property is topologically transitive. Hence $f$ is a topologically transitive Anosov diffeomorphism.
$(6)\Rightarrow(7)$ Suppose that $f$ is a topologically transitive Anosov diffeomorphism. So $f$ is an expansive topologically transitive diffeomorphism that has the shadowing property. Since $M$ is connected, then $f$ is topologically mixing [15, Proposition 3]. Therefore, by [30, Theorem 45], $f$ has the specification property.
$(7)\Rightarrow(8)$ This is a consequence of [47, Proposition 2.1]. See also the remarks on [60, page 4] and [29, page 249].
$(8)\Rightarrow(9)$ This is [29, Theorem 3.5].
$(9)\Rightarrow(10)$ This is [29, Theorem 3.7].
$(10)\Rightarrow(11)$ This is [29, Lemma 3.1].
$(11)\Rightarrow(1)$ Since $f$ is $\Omega-$stable and, in addition, chain-transitive, we conclude as previously that $\Omega(f)=\overline{\operatorname{Per}(f)}=CR(f)=M$ and $\Omega(f)$ is hyperbolic. So $f$ is a chain-transitive diffeomorphism satisfying the shadowing property, and therefore topologically transitive. By [14, Corollary 2.3], the topologically transitive Anosov diffeomorphism $f$ has the two-sided limit shadowing property.
@@
As explained before, by the $\Omega-$stability of $f$, one has $\Omega(f)=\overline{\operatorname{Per}(f)}=CR(f)$ and $CR(f)$ is hyperbolic. Besides, by definition, every periodic point of $f$ is an element of the measure center $\mathcal{M}(M,f)$. Moreover, due to [16, Lemma 5.2], one has
@@
Now, with the equalities $\overline{\mathcal{P}_{f}^{erg}(M)}=\mathcal{P}_{f}(M)$ and $\mathcal{M}(M,f)=CR(f)$, we deduce from [16, Lemma 5.8] that $f$ is topologically transitive. In particular, $f$ is chain-transitive.
@@
Applying Lemma 4.2 when $\xi=\delta$, we deduce that there exists $N=N(\delta)$ such that there are:
@@
  • a $\delta-$chain $(x_{k-1,k})_{0}^{N-1}$ from $T^{n_{k-1}}(y_{k-1})$ to $y_{k}$ with length $p_{(k-1)}$, where
    $(x_{k-1,k})_{0}\,=\,T^{n_{k-1}}(y_{k-1})\quad\quad\text{and}\quad\quad(x_{k-1,k})_{p_{(k-1)}-1}\,=\,y_{k}$
    and $2\leq p_{(k-1)}\leq N$.
@@
In [16, Proposition 5.7], it was proved that, assuming the shadowing property, then the topological transitivity is equivalent to the denseness of the ergodic measures in the space of invariant probability measures.
@@
The proof of Theorem C also ensures, without the assumption of $\Omega-$stability, that
two-sided limit shadowing $\Rightarrow$ $su-$intersecting property in $M$. (6.1)
Moreover (see [53, Lemma 2.1]),
specification propertyweak suintersecting property in $\text{specification property}\quad\Rightarrow\quad\text{weak $su-$intersectingproperty in $\operatorname{Per}_{h}(f)$}.$
@@ 7. Applications
Let $M$ be a smooth closed Riemannian manifold and ${\rm Diff}^{1}(M)$ stand for the set of all $C^{1}$ diffeomorphisms of $M$, endowed with the $C^{1}$ topology. On the next pages we discuss a few consequences of our main results. They comprise the description of both the $C^{1}$ interior and a $C^{1}$ generic subset of the space of diffeomorphisms which satisfy the barycenter property in the set of periodic points; and the proof that, $C^{1}$ generically in the space of volume-preserving diffeomorphisms, to satisfy the barycenter property in the set of hyperbolic periodic points is equivalent to be a transitive Anosov diffeomorphism.
@@ 7.1. Application 1
There is a $C^{1}$ Baire generic set $\mathcal{R}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}$ and there exists $A\subseteq M$ such that $f(A)=A$, $\Omega(f)=\overline{A}$ and $f$ has the barycenter property in $A$, then $f$ is topologically mixing.
@@
Recall that, from [10, Corollaire 1.2], there exists a Baire generic set $\mathcal{R}_{1}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}_{1}$, then $\Omega(f)=CR(f)$. Moreover, as we are assuming that $M$ is connected, by [10, Corollaire 1.3] there exists a Baire generic set $\mathcal{R}_{2}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}_{2}$ and $M=CR(f)$, then $f$ is topologically transitive. From [3, Theorem 2], there exists a Baire generic set $\mathcal{R}_{3}\subset{\rm Diff}^{1}(M)$ such that, if $f\in\mathcal{R}_{3}$ and $f$ is topogically transitive, then $f$ is topogically mixing. Take $f\in\mathcal{R}=\mathcal{R}_{1}\cap\mathcal{R}_{2}\cap\mathcal{R}_{3}$ and suppose that, for some $A\subseteq M$, we have $f(A)=A$, $\Omega(f)=\overline{A}$ and $f$ has the barycenter property in $A$. Therefore, by Lemma 4.1, $f$ is chain-transitive; and so $CR(f)=M$. Hence, $f$ is topologically transitive. Since $f\in\mathcal{R}_{3}$, then $f$ is topologically mixing. ∎
@@ 7.2. Application 2
Since an Anosov diffeomorphism $f\colon M\to M$ is structurally stable (so the intersections of every stable manifold with every unstable manifold are transversal) and, by Theorem C, a topologically transitive Anosov diffeomorphism has the weak $su-$intersecting property in $\operatorname{Per}(f)$, we conclude that, under the transitivity assumption, every two periodic points of $f$ are homoclinically related. Consequently, by Theorem C we conclude that:
@@
Adapting the proof of [14, Corollary 3.2] we obtain:
@@
Let $p,q$ be two periodic points of $f$. As $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, the sets $W^{s}_{f}(\mathcal{O}_{f}(p))\cap W^{u}_{f}(\mathcal{O}_{f}(q))$ and $W^{u}_{f}(\mathcal{O}_{f}(p))\cap W^{s}_{f}(\mathcal{O}_{f}(q))$ are nonempty. Taking into account that $f\in\mathcal{KS}$, we know that these intersections are transversal. Therefore, $p$ and $q$ are homoclinically related, and so they have the same index. ∎
@@
$\mathcal{PCI}\,=\,\big{\{}f\in{\rm Diff}^{1}(M)\colon\,\forall\,p,q\in\operatorname{Per}_{h}(f),\,\operatorname{ind}(p)=\operatorname{ind}(q)\big{\}}.$
@@
[53, Lemma 2.4] Let $G\in{\rm Diff}^{1}(M)$ and $P$ be a non-hyperbolic periodic point of $G$. Then, for every $C^{1}$ neighborhood $\mathcal{V}$ of $G$ there are $H\in\mathcal{V}$ and $P_{1},P_{2}\in\operatorname{Per}_{h}(H)$ such that $\operatorname{ind}(P_{1})\neq\operatorname{ind}(P_{2})$.
By Lemma 7.5, there exist $h\in\mathcal{U}$ and two distinct hyperbolic periodic points $p_{1},p_{2}$ of $h$ with different indices. Yet, this contradicts the fact that $h\in\mathcal{PCI}$. Therefore, $f$ is a star diffeomorphism, and so $\Omega-$stable. ∎
@@
Suppose that $f$ is chain-transitive and $f$ is in $C^{1}$ interior of $\mathcal{PCI}$. By Proposition 7.4, the diffeomorphism $f$ is $\Omega-$stable. Since $f$ is also chain-transitive, by Theorem C $f$ is a topologically transitive Anosov diffeomorphism. The converse assertion is clear, due to the openess of the set of topologically transitive Anosov diffeomorphisms and Corollary 7.2. ∎
The next proposition complements the previous information on a $C^{1}$ generic subset of ${\rm Diff}^{1}(M)$.
@@
[14, Lemma 3.3] There exists an $C^{1}$ open and dense set $\mathcal{I}$ of ${\rm Diff}^{1}(M)$ such that every $f\in\mathcal{I}$ has this property: if for any $C^{1}$ neighborhood $\mathcal{U}$ of $f$ some $g\in\mathcal{U}$ has two hyperbolic periodic points with different indices, then $f$ also has two hyperbolic periodic points with different indices.
@@
If $f\in\mathcal{R}\,\cap\,\mathcal{PCI}$ and is chain-transitive, then $f$ is a topologically transitive Anosov diffeomorphism.
@@
Take a chain-transitive $f\in\mathcal{R}\,\cap\,\mathcal{PCI}$. By Proposition 7.7, $f$ is $\Omega-$stable. Then, by Theorem C, $f$ is a topologically transitive Anosov diffeomorphism. ∎